OpenStudy (anonymous):

Solve the inequality: -1 + 4y < or equal to 31?

OpenStudy (ccswims):

add 1 to 31 first

OpenStudy (ccswims):

You should get \[4y \le 32\]

OpenStudy (ccswims):

now cancel out 4 by division, So technically our next step if 32/4 or 32 divided by 4

OpenStudy (ccswims):

Can you do the division yourself?

OpenStudy (anonymous):

i got y< equal to 8

OpenStudy (anonymous):

you solve inequalities the same way you solve equations...except with inequalities, if you divide/multiply by a negative, the sign changes...but that does not apply in this problem. -1 + 4y <= 31 --- add 1 to both sides 4y <= 31 + 1 4y <= 32 -- now divide both sides by 4 y <= 32/4 y <= 8

OpenStudy (ccswims):

yes

OpenStudy (anonymous):

thank you guys!

OpenStudy (anonymous):

thats what we are here for :)